# Online Posterior Sampling with a Diffusion Prior

### v1

@@ A.2 Proof of Theorem 2
Derivation of $p(s_{t-1}\mid s_{t},h)$. From $p(h\mid s_{0})\propto\mathcal{N}(s_{0};\bar{\theta},\bar{\Sigma})$ and (6), it immediately follows that
$\int_{s_{0}}p(h\mid s_{0})\,p(s_{0}\mid s_{t-1})\mathop{}\!\mathrm{d}s_{0}\approx\mathcal{N}(s_{t-1}/\sqrt{\bar{\alpha}_{t-1}};\bar{\theta},\bar{\Sigma})\propto\mathcal{N}(s_{t-1};\sqrt{\bar{\alpha}_{t-1}}\bar{\theta},\bar{\alpha}_{t-1}\bar{\Sigma})\,.$
The last step treats $\bar{\alpha}_{t-1}$ and $\bar{\Sigma}$ as constants, because the forward process, $t$, and evidence are fixed. Now we apply Lemma 6 to distributions
@@
$p(s_{t-1}\mid s_{t},h)\approx\mathcal{N}(s_{t-1};\hat{\mu}_{t}(s_{t},h),\hat{\Sigma}_{t}(h))\,,$
where $\hat{\mu}_{t}(s_{t},h)$ and $\hat{\Sigma}_{t}(h)$ are defined in the claim. This distribution is a product of two Gaussians: the prior with mean $\mu_{t}(s_{t})$ and covariance $\Sigma_{t}$, and the evidence with mean $\bar{\theta}$ and covariance $\bar{\Sigma}$.
@@
$\int_{s_{0}}p(h\mid s_{0})\,p(s_{0}\mid s_{T})\mathop{}\!\mathrm{d}s_{0}\approx\mathcal{N}(s_{T};\sqrt{\bar{\alpha}_{T}}\bar{\theta},\bar{\alpha}_{T}\bar{\Sigma})\,.$
Then we apply Lemma 6 to distributions
@@
$p(s_{T}\mid h)\approx\mathcal{N}(s_{T};\hat{\mu}_{T+1}(h),\hat{\Sigma}_{T+1}(h))\,,$
where $\hat{\mu}_{T+1}(h)$ and $\hat{\Sigma}_{T+1}(h)$ are given in the claim. This distribution is a product of two Gaussians: the prior with mean $\mathbf{0}_{d}$ and covariance $I_{d}$, and the evidence with mean $\bar{\theta}$ and covariance $\bar{\Sigma}$.
A.3 Proof of Theorem 3
@@
where we introduce $\bar{\theta}$ from Section 2.1. Now we bound each term on the right-hand side.
Upper bound on $\|\tilde{\theta}-\bar{\theta}\|_{2}$. This part of the proof is based on analyzing the asymptotic behavior of the conditional densities in Theorem 2.
@@
[…]
@@
Moreover, $\lambda_{d}(\bar{\Sigma}^{-1})\to\infty$ implies $\lambda_{1}(\bar{\Sigma})\to 0$, and thus $\lim_{N\to\infty}\|S_{t-1}-\sqrt{\bar{\alpha}_{t-1}}\bar{\theta}\|_{2}=0$.
In the last stage, $t=1$, $\bar{\alpha}_{0}=1$, and $S_{0}=\tilde{\theta}$, which implies that
$\lim_{N\to\infty}\|\tilde{\theta}-\bar{\theta}\|_{2}\to 0\,.$
Upper bound on $\|\bar{\theta}-\theta_{*}\|_{2}$. This part of the proof uses the definition of $\bar{\theta}$ in Section 2.1 and that $\varepsilon_{\ell}\sim\mathcal{N}(0,\sigma^{2})$ is independent noise. By definition,
@@ A.4 Proof of Theorem 4
$\int_{s_{0}}p(h\mid s_{0})\,p(s_{0}\mid s_{t-1})\mathop{}\!\mathrm{d}s_{0}\approx p(h\mid s_{t-1}/\sqrt{\bar{\alpha}_{t-1}})\,.$
@@
$p(h\mid\gamma s_{t-1})\,\mathcal{N}(\gamma s_{t-1};\gamma\mu_{t}(s_{t}),\gamma^{2}\Sigma_{t})=\mathcal{N}(\gamma s_{t-1};\dot{\theta}_{t},\dot{\Sigma}_{t})\propto\mathcal{N}(s_{t-1};\dot{\theta}_{t}/\gamma,\dot{\Sigma}_{t}/\gamma^{2})\,,$
where $\dot{\theta}_{t},\dot{\Sigma}_{t}\leftarrow{\tt IRLS}(\gamma\mu_{t}(s_{t}),\gamma^{2}\Sigma_{t},h)$.
@@
$\int_{s_{0}}p(h\mid s_{0})\,p(s_{0}\mid s_{T})\mathop{}\!\mathrm{d}s_{0}\approx p(h\mid s_{T}/\sqrt{\bar{\alpha}_{T}})\,.$
@@
$p(h\mid\gamma s_{t-1})\,\mathcal{N}(s_{t-1};\mathbf{0}_{d},I_{d})\propto\mathcal{N}(s_{t-1};\dot{\theta}_{T+1}/\gamma,\dot{\Sigma}_{T+1}/\gamma^{2})$
where $\dot{\theta}_{T+1},\dot{\Sigma}_{T+1}\leftarrow{\tt IRLS}(\mathbf{0}_{d},\gamma^{2}I_{d},h)$.
A.5 Supplementary Lemmas
@@
The neglected factors depend on $\mu_{1}$, $\mu_{2}$, $\Sigma_{1}$, and $\Sigma_{2}$. This completes the proof. ∎
@@ Appendix B Learning the Reverse Process
One property of our model is that $q(s_{T})\approx\mathcal{N}(s_{T};\mathbf{0}_{d},I_{d})$ when $T$ is sufficiently large [18]. Since $S_{T}$ is initialized to the same distribution in the reverse process $p$, $p$ can be learned from the forward process $q$ by simply reversing it. This is done as follows. Using the definition of the forward process in (4), Ho et al. [18] showed that
@@
Therefore, the latent variable in stage $t-1$, $S_{t-1}$, is easy to sample when $s_{t}$ and $s_{0}$ are known. To estimate $s_{0}$, which is unknown when sampling from the reverse process, we use the forward process again. In particular, (4) implies that $s_{t}=\sqrt{\bar{\alpha}_{t}}s_{0}+\sqrt{1-\bar{\alpha}_{t}}\tilde{\varepsilon}_{t}$ for any $s_{0}$, where $\tilde{\varepsilon}_{t}\sim\mathcal{N}(\mathbf{0}_{d},I_{d})$ is a standard Gaussian noise. This identity can be rearranged as
$s_{0}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}(s_{t}-\sqrt{1-\bar{\alpha}_{t}}\tilde{\varepsilon}_{t})\,,$
To obtain $\tilde{\varepsilon}_{t}$, which is unknown when sampling from $p$, we learn to regress it from $s_{t}$ [18].
The regressor is learned as follows. Let $\varepsilon_{t}(\cdot;\psi)$ be a regressor of $\tilde{\varepsilon}_{t}$ parameterized by $\psi$ and $\mathcal{D}=\left\{s_{0}\right\}$ be a dataset of training examples. We sample $s_{0}$ uniformly at random from $\mathcal{D}$ and then solve
$\psi_{t}=\operatorname*{arg\,min\,}_{\psi}\mathbb{E}_{q}\left[\|\tilde{\varepsilon}_{t}-\varepsilon_{t}(S_{t};\psi)\|_{2}^{2}\right]$ (14)
per stage. The expectation is approximated by an empirical average over sampled $s_{0}$. Note that we slightly depart from Ho et al. [18]. Since each $\varepsilon_{t}$ has its own parameters, the original optimization problem over $T$ stages decomposes into $T$ subproblems.
@@ Appendix C Additional Experiments
This section contains three additional experiments.
@@ C.1 Additional Synthetic Problems
Figure 4: Evaluation of $\tt DiffTS$ on another three synthetic problems. The first row shows samples from the true (blue) and diffusion model (red) priors. The second row shows the regret of $\tt DiffTS$ and the baselines as a function of round $n$.
In Section 6.2, we show results for three hand-selected problems out of six. We report results on the other three problems in Figure 4. We observe the same trends as in Section 6.2. The only minor change is that $\tt DPS$ does not diverge on the last problem.
@@ C.2 MNIST Experiment
Figure 5: Evaluation of $\tt DiffTS$ on the MNIST dataset.
The next experiment is on the MNIST dataset [30], where the mean reward is estimated by a classifier. We start with learning an MLP-based multi-way classifier for digits and extract $d=8$ dimensional embeddings of all digits in the dataset, which are used as feature vectors in our experiment. We generate a distribution over model parameters $\theta_{*}$ as follows: (1) we choose a random positive label, assign it reward $1$, and assign reward $-1$ to all other labels; (2) we subsample a random dataset of size $20$, with $50\%$ positive and $50\%$ negative labels; (3) we train a linear model, which gives us a single $\theta_{*}$. We repeat this 10 0001000010\,00010 000 times and get a distribution over $\theta_{*}$.
We consider both linear and logistic bandits. In both, the model parameter $\theta_{*}$ is initially sampled from the prior. In each round, $K=10$ random actions are chosen randomly from all digits. In the linear bandit, the mean reward for digit $x$ is $x^{\top}\theta_{*}$ and the reward noise is $\sigma=1$. In the logistic bandit, the mean reward is $g(x^{\top}\theta_{*})$ where $g$ is a sigmoid.
Our MNIST results are reported in Figure 5. We observe again that $\tt DiffTS$ has a lower regret than all baselines, because the learned prior captures the underlying distribution of $\theta_{*}$ well. We note that both the prior and diffusion prior distributions exhibit a strong cluster structure (Figure 5a), where each cluster represents one label.
@@ C.3 Ablation Studies
Figure 6: An ablation study of $\tt DiffTS$ on the cross problem: (a) regret with a varying number of samples for training the diffusion prior, (b) regret with a varying number of diffusion stages $T$, and (c) computation time with a varying number of diffusion stages $T$.
We conduct three ablation studies on the cross problem in Figure 2.
In all experiments, the number of samples for training diffusion priors was 10 0001000010\,00010 000. In Figure 6a, we vary it from $100$ to 10 0001000010\,00010 000. We observe that the regret decreases as the number of samples increases, due to learning a more accurate prior approximation. It stabilizes around 3 00030003\,0003 000 training samples. We conclude that the quality of the learned prior approximation has a major impact on regret.
In all experiments, the number of diffusion stages was $T=100$. In Figure 6b, we vary it from $1$ to $300$ and observe its impact on regret. The regret at $T=1$ is high, but then it drops quickly up to $T=25$. It stabilizes around $T=100$, which we used in our experiments. In Figure 6c, we vary $T$ from $1$ to $300$ and observe its impact on the computation time of posterior sampling. As suggested in Section 4.2, it increases linearly with $T$. The main contribution to the computation time are $T$ calls to the neural network regressor.

### v2

@@ A.2 Proof of Theorem˜2
Derivation of $p(s_{t-1}\mid s_{t},h)$. From (6) and Assumption˜1, it follows that
$\int_{s_{0}}p(h\mid s_{0})\,p(s_{0}\mid s_{t-1})\mathop{}\!\mathrm{d}s_{0}$ $\propto p(h\mid s_{t-1}/\sqrt{\bar{\alpha}_{t-1}})\propto\mathcal{N}(s_{t-1}/\sqrt{\bar{\alpha}_{t-1}};\bar{\theta},\bar{\Sigma})$
$\propto\mathcal{N}(s_{t-1};\sqrt{\bar{\alpha}_{t-1}}\bar{\theta},\bar{\alpha}_{t-1}\bar{\Sigma})\,.$
The last step treats $\bar{\alpha}_{t-1}$ and $\bar{\Sigma}$ as constants, because the forward process and evidence $h$ are fixed. Now we apply Lemma˜6 to distributions
@@
$p(s_{t-1}\mid s_{t},h)\propto\mathcal{N}(s_{t-1};\hat{\mu}_{t}(s_{t},h),\hat{\Sigma}_{t}(h))\,$,
where $\hat{\mu}_{t}(s_{t},h)$ and $\hat{\Sigma}_{t}(h)$ are defined in the claim. This is a product of two Gaussians: the prior with mean $\mu_{t}(s_{t})$ and covariance $\Sigma_{t}$, and the evidence with mean $\sqrt{\bar{\alpha}_{t-1}}\bar{\theta}$ and covariance $\bar{\alpha}_{t-1}\bar{\Sigma}$.
@@
$\int_{s_{0}}p(h\mid s_{0})\,p(s_{0}\mid s_{T})\mathop{}\!\mathrm{d}s_{0}\propto\mathcal{N}(s_{T};\sqrt{\bar{\alpha}_{T}}\bar{\theta},\bar{\alpha}_{T}\bar{\Sigma})\,.$
Then we apply Lemma˜6 to distributions
@@
$p(s_{T}\mid h)\propto\mathcal{N}(s_{T};\hat{\mu}_{T+1}(h),\hat{\Sigma}_{T+1}(h))\,$,
where $\hat{\mu}_{T+1}(h)$ and $\hat{\Sigma}_{T+1}(h)$ are defined in the claim. This is a product of two Gaussians: the prior with mean $\mathbf{0}_{d}$ and covariance $I_{d}$, and the evidence with mean $\sqrt{\bar{\alpha}_{T}}\bar{\theta}$ and covariance $\bar{\alpha}_{T}\bar{\Sigma}$.
A.3 Proof of Theorem˜3
@@
where we introduce $\bar{\theta}$ from Section˜2.1. Now we bound each term on the right-hand side.
Upper bound on $\|\tilde{\theta}-\bar{\theta}\|_{2}$. This part of the proof is based on analyzing the asymptotic behavior of the conditional densities in Theorem˜2.
@@
$\hat{\mu}_{T+1}(h)=\hat{\Sigma}_{T+1}(h)(I_{d}\,\mathbf{0}_{d}+\bar{\Sigma}^{-1}\bar{\theta}/\sqrt{\bar{\alpha}_{T}})\,,\quad\hat{\Sigma}_{T+1}(h)=(I_{d}+\bar{\Sigma}^{-1}/\bar{\alpha}_{T})^{-1}\,.$
@@
Moreover, $\lambda_{d}(\bar{\Sigma}^{-1})\to\infty$ implies $\lambda_{1}(\bar{\Sigma})\to 0$, and thus $\lim_{N\to\infty}\|S_{t-1}-\sqrt{\bar{\alpha}_{t-1}}\bar{\theta}\|_{2}=0$ for any $t\in[T]$. In the last stage, $t=1$, $\bar{\alpha}_{0}=1$, and $S_{0}=\tilde{\theta}$. Therefore,
$\lim_{N\to\infty}\|S_{t-1}-\sqrt{\bar{\alpha}_{t-1}}\bar{\theta}\|_{2}=\lim_{N\to\infty}\|\tilde{\theta}-\bar{\theta}\|_{2}=0\,.$
Upper bound on $\|\bar{\theta}-\theta_{*}\|_{2}$. This part of the proof uses the definition of $\bar{\theta}$ in Section˜2.1 and that $\varepsilon_{\ell}\sim\mathcal{N}(0,\sigma^{2})$ is independent noise. By definition,
@@ A.4 Proof of Theorem˜4
$\int_{s_{0}}p(h\mid s_{0})\,p(s_{0}\mid s_{t-1})\mathop{}\!\mathrm{d}s_{0}\propto p(h\mid s_{t-1}/\sqrt{\bar{\alpha}_{t-1}})\,.$
@@
$p(h\mid\gamma s_{t-1})\,\mathcal{N}(\gamma s_{t-1};\gamma\mu_{t}(s_{t}),\gamma^{2}\Sigma_{t})\propto\mathcal{N}(\gamma s_{t-1};\dot{\theta}_{t},\dot{\Sigma}_{t})\propto\mathcal{N}(s_{t-1};\dot{\theta}_{t}/\gamma,\dot{\Sigma}_{t}/\gamma^{2})\,$,
where $\dot{\theta}_{t},\dot{\Sigma}_{t}\leftarrow{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt IRLS}(\gamma\mu_{t}(s_{t}),\gamma^{2}\Sigma_{t},h)$.
@@
$\int_{s_{0}}p(h\mid s_{0})\,p(s_{0}\mid s_{T})\mathop{}\!\mathrm{d}s_{0}\propto p(h\mid s_{T}/\sqrt{\bar{\alpha}_{T}})\,.$
@@
$p(h\mid\gamma s_{T})\,\mathcal{N}(s_{T};\mathbf{0}_{d},I_{d})\propto\mathcal{N}(s_{T};\dot{\theta}_{T+1}/\gamma,\dot{\Sigma}_{T+1}/\gamma^{2})\,$,
where $\dot{\theta}_{T+1},\dot{\Sigma}_{T+1}\leftarrow{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt IRLS}(\mathbf{0}_{d},\gamma^{2}I_{d},h)$.
A.5 Supporting Lemmas
@@
The neglected factors depend on constants $\mu_{1}$, $\mu_{2}$, $\Sigma_{1}$, and $\Sigma_{2}$. This completes the proof. ∎
@@ Appendix B Learning the Reverse Process
One property of our model is that $q(s_{T})\approx\mathcal{N}(s_{T};\mathbf{0}_{d},I_{d})$ when $T$ is sufficiently large [19]. Since $S_{T}$ has the same distribution in the reverse process $p$, $p$ can be learned from the forward process $q$ by simply reversing it. This is done as follows. Using the definition of the forward process in (4), Ho et al. [19] showed that
@@
Therefore, the latent variable in stage t1t-1, $S_{t-1}$, is easy to sample when $s_{t}$ and $s_{0}$ are known. To estimate $s_{0}$, which is unknown when sampling from the reverse process, we use the forward process again. In particular, (4) implies that $s_{t}=\sqrt{\bar{\alpha}_{t}}s_{0}+\sqrt{1-\bar{\alpha}_{t}}\varepsilon_{t}$, where $\varepsilon_{t}\sim\mathcal{N}(\mathbf{0}_{d},I_{d})$ is a standard Gaussian noise. This identity can be rearranged as
$s_{0}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}(s_{t}-\sqrt{1-\bar{\alpha}_{t}}\varepsilon_{t})\,.$
To obtain $\varepsilon_{t}$, which is unknown when sampling from $p$, we learn to regress it from $s_{t}$ [19].
The regressor is learned as follows. Let $\varepsilon_{t}(\cdot;\psi)$ be a regressor of $\varepsilon_{t}$ parameterized by $\psi$ and $\mathcal{D}=\left\{s_{0}\right\}$ be a dataset of training examples. We sample $s_{0}$ uniformly at random from $\mathcal{D}$ and then solve
$\psi_{t}=\operatorname*{arg\,min\,}_{\psi}\mathbb{E}_{q}\left[\|\varepsilon_{t}-\varepsilon_{t}(S_{t};\psi)\|_{2}^{2}\right]$ (15)
per stage. The expectation is approximated by sampled $s_{0}$. Note that we slightly depart from Ho et al. [19]. Since each regressor has its own parameters, the original optimization problem over $T$ stages decomposes into $T$ subproblems.
@@ Appendix C Additional Experiments
This section contains four additional experiments.
@@ C.1 Additional Synthetic Problems
Figure 4: Evaluation of $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DiffTS$ on another three synthetic problems. The first row shows samples from the true (blue) and diffusion model (red) priors. The second row shows the regret of $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DiffTS$ and the baselines as a function of round $n$.
In Section˜6.2, we show results for three hand-selected problems out of six. We report results on the other three problems in Figure˜4. We observe the same trends as in Section˜6.2.
@@ C.2 MNIST Experiment
Figure 5: Evaluation of $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DiffTS$ on the MNIST dataset: (a) shows samples from the true (blue) and diffusion model (red) priors, (b) shows regret in the linear bandit, and (c) shows regret in the logistic bandit.
The next experiment is on the MNIST dataset [31]. We start with learning an MLP-based multi-way classifier for digits and extract their $d=8$ dimensional embeddings. These are used as features in our experiment. We generate a distribution over model parameters $\theta_{*}$ as follows: (1) we choose a random positive label, assign it reward $1$, and assign reward 1-1 to all other labels; (2) we subsample a random dataset of size $20$, with $50\%$ positive and $50\%$ negative labels; (3) we train a linear model, which gives us a single $\theta_{*}$. We repeat this 10 00010\,000 times and get a distribution over $\theta_{*}$.
We consider both linear and logistic bandits. In both, the model parameter $\theta_{*}$ is initially sampled from the prior. In each round, $K=10$ random actions are chosen randomly from all digits. In the linear bandit, the mean reward for a digit with embedding $x$ is $x^{\top}\theta_{*}$ and the reward noise is $\sigma=1$. In the logistic bandit, the mean reward is $g(x^{\top}\theta_{*})$, where $g$ is a sigmoid.
Our MNIST results are reported in Figure˜5. We observe again that $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DiffTS$ has a lower regret than all baselines, because the learned prior captures the underlying distribution of $\theta_{*}$ well. We note that both the prior and diffusion prior distributions exhibit a strong cluster structure (Figure˜5a), where each cluster represents one label.
@@ C.3 Ablation Studies
Figure 6: An ablation study of $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DiffTS$ on the cross problem: (a) we vary the number of samples for training the diffusion prior and report regret, (b) we vary the number of diffusion stages $T$ and report regret, and (c) we vary the number of diffusion stages $T$ and report computation time.
Figure 7: Evaluation on Gaussian mixture variants of the synthetic problems in Figure 2. The first row shows samples from the true (blue) and diffusion model (red) priors. The second row shows the earth mover’s distance of $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DiffTS$ and baseline posteriors from the true posterior as a function of sample size $n$.
We conduct three ablation studies on the cross problem in Figure˜2.
In all experiments, the number of samples for training diffusion priors was 10 00010\,000. In Figure˜6a, we vary it from $100$ to 10 00010\,000. We observe that the regret decreases as the number of samples increases, due to learning a better prior approximation. The trend stabilizes around 3 0003\,000 training samples. We conclude that the quality of the learned prior approximation has a major impact on regret.
In all experiments, the number of diffusion stages was $T=100$. In Figure˜6b, we vary it from $1$ to $300$ and observe its impact on regret. While the regret at $T=1$ is high, it decreases quickly as $T$ increases. It stabilizes around $T=100$, which we used in our experiments. In Figure˜6c, we vary $T$ from $1$ to $300$ and observe its effect on the computation time of posterior sampling. The time is linear in $T$, as suggested in Section˜4.2. The main contributor is the neural network regressor.
C.4 Non-Bandit Evaluation
We use Gaussian mixture variants of the synthetic problems in Figure˜2 for our non-bandit evaluation. The action in round $k$ is chosen uniformly at random (not adaptively). Since the priors are Gaussian mixtures, the true posterior distribution can be computed in a closed form using $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt MixTS$ and we can measure the distance of posterior approximations from it. We use the earth mover’s distance (EMD) between posterior samples from the true posterior and its approximation. We also considered the KL divergence. However, we could not apply it because the posteriors of $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DiffTS$ and $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DPS$ do not have analytical forms.
We evaluate all methods from Figure˜2. In addition, we implement a sequential Monte Carlo (SMC) sampler [16]. The initial particles are sampled uniformly at random from the prior. At each round, the particles are perturbed by a Gaussian noise. The standard deviation of the noise is initialized as a fraction of the observation noise and decays over time, as the posterior concentrates. The particles are weighted according to the likelihood of the observation in the round. Finally, we use normalized likelihood weights to resample the particles. The number of particles is 3 0003\,000 and we tune SMC to get good posterior approximations. The computational cost of SMC is comparable to $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DiffTS$.
Our results are reported in Figure˜7. We observe that $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DiffTS$ approximations are comparable to $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt MixTS$, which has an exact posterior in this setting. The second best performing method is SMC. Its approximations worsen as the sample size $n$ increases. $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DPS$ approximations also get worse as $n$ increases, which caused instability in Figure˜2.